INAF queue observing, 2015 Nov 05-Nov 06

Observer: S. Bisogni, R. Carini
Telescope Operator: Steve Allanson
Support Astronomer: Barry Rothberg
Instrument: LUCI

Summary

22:40 UT the humidity is still high (95%), but the weather is really sunny. The telescope operator is going to check the roof to see if we can can get rid of the melted snow on it and open, but he's still not sure about the humidity situation..
00:00 still have a lot of rime ice on the weather masts, lightning rods, railings, etc. Some ice remains on the shutter seals. Falling ice has been pounding the lounge roof and high bay all day. The TO, John Little, Jenny, Barry, and Christian agree that this is an unsafe situation in which to open. Furthermore, we are steady at ~95% RH. The rear anemometer is frozen and unresponsive.
The dome is closed

Night duration (12\xB0 el twilight) 19:00-5:10 ( 02:00-12:10 UT) = 10 h 10m

Observing time: 0 hrs 0 m

Weather time loss: 10 h 10 m (100%)

technical time loss 0 m

Data Summary

Project_RUN Objects Filter ExpTime Notes
Iovino calibrazione HK, zJ      
  Bias LBC      

Details

04:25 test for flats and lamps for Iovino's program luci2.20151106.0001-0003

05:15 GO! 2015_2016_5_HK_spec_new_cal.xml luci2.20151106-0004-0031

05:36 GO! 2015_2016_5_zJ_spec_new_cal.xml luci2.20151106-0032-0059

08:49 bias with LBC: Run 25BIAS_BINO

Article text.

-- %USERSIG{SusannaBisogni - 2015-11-05}%

Comments

 
Topic revision: r6 - 07 Nov 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback